OpenStudy (anonymous):

PLEASE HELP! Write each of the following in terms of sin θ and cos θ, and then simplify if possible. csc θ cot θ

OpenStudy (lgbasallote):

hint: \[\csc \theta = \frac{1}{\sin \theta}\] \[\cot \theta = \frac{\cos \theta}{\sin \theta}\]

OpenStudy (anonymous):

so I multiply

OpenStudy (nali):

still need help?

OpenStudy (anonymous):

yes

OpenStudy (nali):

it is just as u said multiply

OpenStudy (nali):

did u get it ?

OpenStudy (anonymous):

so the answer is cos θ/ sin^2 θ

OpenStudy (nali):

yes that is right

OpenStudy (anonymous):

ok thanks

OpenStudy (nali):

ur welcome